### v1

# SparsePixels: Efficient Convolution for Sparse Data on FPGAs

Abstract.
Inference of standard convolutional neural networks (CNNs) on Field-Programmable Gate Arrays (FPGAs) often incurs high latency and a long initiation interval due to the deep nested loops required to densely convolve every input pixel regardless of its feature value, especially when the image size is large. However, in some image data, input features can be spatially sparse, and semantic information may occupy only a small fraction of the input pixels. In this case most computation would be wasted on empty regions. In this work, we introduce SparsePixels, a framework for efficient convolution for spatially sparse image data on FPGAs, targeting fast inference applications in constrained environments with latency requirements of microseconds or below. Our approach implements a special class of CNNs that selectively retain and compute on a small subset of pixels that are active while ignoring the rest. We show that, for example, in a neutrino physics dataset for identifying neutrino interactions in LArTPC images that have around 4k input pixels but are naturally very sparse, a standard CNN with a compact size of 4k parameters incurs an inference latency of 48.665 s on an FPGA, whereas a sparse CNN of the same base architecture computing on less than 1% of the input pixels results in a inference speedup to 0.665 s, with resource utilization well within on-chip budgets, trading only a small percent-level performance loss. At least one-order-of magnitude speedups with comparable performance are also demonstrated in similar datasets with sparse image patterns. This work aims to benefit future algorithm developments for fast and efficient data readout in modern experiments such as the trigger and data acquisition systems at the CERN Large Hadron Collider. For easy adoption, we have developed a library to support building sparse CNNs with quantization-aware training, as well as an HLS implementation for FPGA deployment. Codes are available at https://github.com/hftsoi/sparse-pixels.
1. Introduction
Convolutional neural networks (CNNs) have gained wide interest and usage for image-based recognition tasks in particle physics experiments, including muon tracking in gas chambers (Francescato et al., 2021; Coccaro et al., 2023; Maglianella et al., 2023), jet flavor tagging in calorimeters (de Oliveira et al., 2016; Kasieczka et al., 2017; Butter et al., 2019; Moreno et al., 2020; Pol et al., 2021), anomaly detection for new physics searches at colliders (Kasieczka et al., 2021; Aarrestad et al., 2022; Govorkova et al., 2022; Collaboration, 2023, 2024; Gandrakota, 2025), and neutrino identification in liquid argon time projection chambers (LArTPCs) (Aurisano et al., 2016; Acciarri et al., 2017a; Dominé and Terao, 2020; Abi et al., 2020; Chung et al., 2025). This is because often experimental data are encoded in geometrical formats originating from detector layouts and can be represented in fixed-grid views. Hence, it is natural to use CNNs to process and extract signal from this type of data. However, some of these images are spatially very sparse because certain signals leave small traces within large detector volumes. For example, in the dateset presented in (Maglianella et al., 2023), a muon traversing gas chambers leaves a track in a 2D image covering the full muon station, but a single muon track may occupy only 10 out of pixels ( of the input). Similarly, in neutrino experiments, particle tracks in large LArTPC volumes can have, on average, only 0.01% nonzero pixels (Dominé and Terao, 2020). Because these signatures can have irregular shapes and appear anywhere in the image, CNNs must scan the entire input to capture them. Implementing large standard CNNs can therefore incur high inference latency even on high-speed custom processors such as FPGAs (Aarrestad et al., 2021), posing a challenge for data readout in modern experiments that enforce ultra-low latency.
In modern particle experiments, data processing is typically divided into two domains: online (low-latency) processing and offline analysis. Taking the CERN Large Hadron Collider (LHC) as an example (Aad et al., 2020; Collaboration, 2018; Sirunyan et al., 2020; Collaboration, 2020), proton bunches collide at 40 MHz, and each collision is an event. The detector acts like a high-speed camera, recording trajectories and energy deposits across subdetectors. The resulting raw information is about 1 MB per event, leading to a raw data rate of TB/s. Storing all data is impossible due to limited bandwidth and storage capacity, hence a fast filtering system (the trigger) is required to select potentially interesting events for permanent storage in real time, concurrent with the 40 MHz collisions. Because the input streams in at 40 MHz, the Level-1 trigger must decide whether to store an event within a few microseconds, set by buffer capacity, motivating the use of FPGAs/ASICs for real-time processing. It is crucial and challenging to design trigger algorithms for online use that are sensitive enough to select interesting events (as unselected events are discarded permanently) while remaining compact enough to meet low-latency constraints. With the forthcoming High-Luminosity LHC upgrade (Zurbano Fernandez et al., 2020), developing more sophisticated and hardware-efficient online algorithms is essential for handling unprecedented rates and bandwidth. After trigger selections, data are stored and analyzed offline without strict latency limits.
In this paper, we focus on the low-latency domain, targeting inference latencies on the order of microseconds and below. The general workflow for deploying machine learning (ML) models on FPGAs in this regime is well established in the community. The library (FastML Team, 2025; Duarte et al., 2018) is widely used due to its flexible hardware-aware configuration, which supports translating trained models from common Python libraries into High-level Synthesis (HLS) firmware for FPGA deployment. CNNs are supported and have been benchmarked (Aarrestad et al., 2021; Tarafdar et al., 2022; Ghielmetti et al., 2022). For large images (e.g., with or more pixels), inputs typically cannot be fully partitioned to enable parallel processing on an FPGA, so designs fall back to a streaming approach, processing one pixel at a time. While this keeps resource usage small, it often yields high latencies. When the image is sparse, say only a few percent or fewer pixels are nonzero on average, the standard convolution becomes highly inefficient because it must be computed on many empty regions. Therefore, we investigate a class of CNNs that adaptively select and compute only on active pixels, targeting large images with extreme sparsity.
Problem statement.
Can we improve the processing efficiency of CNNs for spatially sparse images on FPGAs to allow for accelerated real-time inference in experiments that require microsecond-level latency or below?
Our contributions.
Development of a sparse convolution method designed for low-latency FPGA implementation.
Proof-of-concept on spatially sparse physics image data, demonstrating significant FPGA inference speedups with comparable model performance.
A Python library for quantization-aware training of sparse CNNs and C++ HLS modules for FPGA deployment.
2. Related work
Our work aims to improve upon the current implementation of standard 2D convolution in (Aarrestad et al., 2021) designed specifically for sparse image data on FPGAs in the low-latency regime, using a sparse convolution method inspired by works on acceleration of sparse 3D point cloud inference.
Low-latency standard convolution on FPGAs. The (Duarte et al., 2018; FastML Team, 2025) library is open-sourced and developed to facilitate deployment of ML models on FPGAs for fast inference in strictly constrained environments such as particle physics experiments at the LHC. The library provides a Python interface for translating models trained in common libraries such as (Chollet et al., 2015; Coelho et al., 2021) and (Paszke et al., 2019) into high-level C/C++ via HLS tools, which can be synthesized into register-transfer level (RTL) designs, enabling rapid customization and optimization of resources and latency for FPGA firmware. has integrated quantization-aware training (QAT) for fixed-precision deployment (Coelho et al., 2021; Sun et al., 2024). Implementation of various ML architectures have been demonstrated (Duarte et al., 2018; Summers et al., 2020; Khoda et al., 2023; Tsoi et al., 2024, 2025; Jiang et al., 2025), including CNNs (Aarrestad et al., 2021; Tarafdar et al., 2022; Ghielmetti et al., 2022), which serve as the baseline in this paper. Other tools such as (Umuroglu et al., 2017; Blott et al., 2018) and (Vreča and Biasizzo, 2025), with slightly different optimization targets and workflows, also support low-latency inference of standard convolution on FPGAs. However, across all these tools, the underlying challenge is similar: the amount of computation required by standard convolution on large inputs makes full parallel processing prohibitive, forcing designs toward streaming implementations (inputs are processed serially) that can incur too high a latency for real-time systems. For example, has demonstrated an inference latency of 5 s for CNNs on images (Aarrestad et al., 2021), and this latency grows rapidly with input size.
Efficient convolution on GPUs. A direction to reduce CNN computation is to exploit sparsity to skip unnecessary operations. (Liu et al., 2015) explored model compression via parameter pruning, reducing redundancy in the parameter space and mapping convolutions to sparse matrix multiplications for speedup. Other works (Lu et al., 2019; Zhu et al., 2020) explored efficient data-flow schemes on FPGAs for CNNs with sparse structure to skip multiplications involving zeros. These methods focus on weight sparsity or structured sparsity, and still operate on all input pixels. Moreover, directly skipping computation based on irregular input sparsity often leads to input-dependent computations and varying runtimes, which is problematic in systems with strict latency bounds. A more relevant direction is to focus on selective computation only on a subset of pixels from the input tensor and ignore the rest. (Graham, 2015) proposed a type of sparse convolution that computes an output pixel only when any of the input pixels within the kernel receptive field is nonzero, so the computational cost scales with the number of nonzero input pixels rather than the entire image/volume. However, this approach can rapidly increase the number of nonzero pixels across convolutional layers: an output pixel can become nonzero if its receptive field includes any neighboring nonzero input, effectively dilating the nonzero set. To address this, (Graham et al., 2018) constrains the operation so that an output pixel is computed only when the same location at the input is nonzero, keeping the sparsity pattern fixed across layers and trading little or no performance loss for significant speedups. (Choy et al., 2019), (Yan et al., 2018), (Tang et al., 2022), and (Tang et al., 2023) are some 3D point cloud engines built upon such sparse convolution for high-performance inference on GPUs.
Despite these advances, such techniques have not yet been explored in the FPGA domain, which is critical for certain scientific experiments that demand especially low latencies. Our proposed method addresses this with an implementation of sparse convolution that selectively computes only on a small subset of input pixels and achieves constant runtime by imposing a maximum output sparsity, enabling parallel processing on FPGAs to achieve large inference speedups.
3. Method
We first present the background and define sparse convolution in Sec. 3.1, then lay out the algorithmic details of the HLS implementation in Sec. 3.2.
3.1. Background and sparse convolution
For simplicity, same padding and unit stride are assumed in the rest of the paper. In a standard 2D convolutional layer, for an input tensor with dimensions (, , ), an output tensor with dimensions (, , ), and a kernel weight tensor with dimensions (, , , ), three nested loops are required to accumulate input-weight multiplications for each output pixel and channel (LeCun et al., 2015):
where denotes the bias terms, and the input indices are shifted to the proper spatial patch. Eq. 1 is computed for each output pixel (, ) and channel (). Thus, a standard Conv2D performs inner products, each with multiplications, totaling multiplications (six nested loops). Directly implementing these deep nested loops in HLS leads to high latency because each transition across loop levels, say exiting an inner loop to its enclosing outer loop or re-entering for the next iteration, costs one clock cycle, introducing loop-carried dependencies and scheduling overhead. A common workaround is to pipeline an outer loop and unroll inner loops so iterations execute in parallel, but this is infeasible for typical image sizes due to resource usage and memory bandwidth. Because of these limitations, uses stream I/O to implement convolutional architectures. In stream I/O, inputs are processed sequentially via “first in, first out” (FIFO) buffers, which use fewer resources than parallel I/O. For an input of size , the pixels are fed one-by-one, each carrying features. Pixels are buffered until a full sliding window is available, then the multiply-accumulate is performed. Since streaming reads one pixel per clock, the latency is at least cycles to scan the image, plus additional cycles to fill and drain the sliding window, which scale with and . This approach enables large inputs but yields too high latencies for constrained systems.
For large images with spatially sparse features, implementing standard convolution is often prohibitive, or, if feasible, wastes most computation on empty regions that contain little or no semantic information and thus contribute little to feature extraction. In contrast, sparse convolution is a special (constrained) type of convolution that selectively computes on a subset of pixels instead of densely convolving the entire image. Which pixels are selected can depend on the dataset or task, but in general, selection is based on pixel activity. We consider an input pixel active if its feature value exceeds a predefined threshold (often zero or a small noise value). Inspired by (Graham et al., 2018), we constrain the operation so that: an output pixel is active if and only if the input at the same pixel location is active, and only active pixels are computed. This preserves the sparsity pattern between input and output and prevents dilation, as illustrated in Fig. 1.
Building on this sparsity-preserving convolution, we introduce the framework, which consists of five core sparse layers sufficient to build a sparse CNN: (1) sparse input reduction, (2) sparse convolution, (3) sparse activation, (4) sparse pooling, and (5) sparse flattening, illustrated in Fig. 2. The framework is designed around implementing CNNs with sparse convolution in HLS. The sparse input reduction is the first layer after the input, which extracts up to a fixed number of active pixels from the input images and groups them for subsequent processing. The sparse convolutional layer performs a sparsity-preserving convolution that fixes the input-output sparsity. The other sparse layers retain their usual functionality but operate on the stored active pixels. For training, we have developed a Python library to build and train sparse CNNs that mirror the HLS behavior. The training library uses as a backend for QAT (Coelho et al., 2021), and will support high-granularity quantization (HGQ) (Sun et al., 2024) in the future. The HLS implementation of the sparse layers is detailed in the following section.
3.2. HLS implementation
In HLS, the core idea is to store only the active pixels in compact arrays and perform all subsequent computation on these arrays.
3.2.1. Sparse input reduction
The first step is to read out the active pixels from the input, as illustrated in Fig. 3. For an input image with pixels and channels, the active pixels are identified as those whose (channel-0 by convention) feature value exceeds a predefined threshold, and are extracted into two compact sparse arrays: the feature array (storing all channel values for each retained pixel) and the hash array (storing the corresponding height and width coordinates). We fix a maximum number of active pixels per model, , which sets the array sizes: has elements and has elements. Thus, instead of computing over all pixels as in standard convolution, we operate on arrays whose size scales with for sparse data. Unlike in Python, which can allow a variable number of pixels per example, the number of active pixels to be computed must be bounded in HLS because the design is fixed once synthesized into firmware. A naive implementation, say scanning all input pixels once and “pushing” active pixels into an output array via an “if-else”, creates data-dependent write addresses, where the next write index depends on how many active pixels were seen. In HLS this induces loop-carried dependencies that can inflate scheduling and synthesis cost even when the output array is small.
To avoid data-dependent writes, we use a recursive binary reduction scheme to read out the active pixels, as illustrated in Fig. 4. Let and flatten the first channel into a length- array. We then split it into two partitions where the left partition has size (largest power of two less than ) and the right holds the remainder. The split recurses until we reach pairs. The reduction tree depth is therefore
Using a simple pairwise combiner that returns the left element if it is above threshold and otherwise the right, we reduce each level until one element remains, which is the leftmost active pixel from the flat input array. We write this pixel’s feature(s) and coordinates into the first slots of and , set its value to zero in the input array, and repeat the reduction to obtain the next leftmost active. This continues until we have filled up to entries. By convention, the ordering is row-major (top-left to bottom-right). This approach is hardware-efficient as it replaces the data-dependent “push-back” with a fixed-shape reduction tree. The algorithmic details are given in Alg. 1.
3.2.2. Sparse convolution
Next, we perform sparse convolution that takes as inputs the two sparse arrays, as illustrated in Fig. 5. Because the operation preserves the sparsity pattern, the active pixel locations are identical at input and output. Consequently, only the feature array is updated while the hash array remains unchanged.
Since only active pixels can contribute and the input and output share the same active locations, no information from inactive pixels is required. On the sparse arrays, we loop over output pixels; for each output pixel we loop over input pixels and compute the offsets between their coordinates from the hash array. If lies within the receptive field, we map that offset to the corresponding kernel position and accumulate a channel-wise dot product between the input features and the appropriate weights; otherwise, the pair contributes nothing. By construction, if an output location is inactive (padded), it will never get activated in subsequent layers.
For each output channel, the sparse convolution involves two nested loops over active pixels (outer: output active pixels; inner: input active pixels). Because multiplications are based on an offset check and weight lookup, we do not loop over all kernel positions explicitly, hence the iteration count does not scale with the kernel size. This is advantageous for sparse data, since active pixels can be far apart, and increasing to allow long-range interactions does not increase the number of loop iterations. The relative multiply-accumulate cost between standard and sparse convolutions can be seen as:
whenever . The algorithmic details are given in Alg. 2.
3.2.3. Sparse activation
As in standard activation, sparse activation applies an element-wise nonlinearity to the sparse feature array while keeping the hash array unchanged. The number of elements processed is greatly reduced:
whenever . The algorithmic details are given in Alg. 3.
3.2.4. Sparse pooling
For sparse pooling, we first map every retained pixel’s coordinate in the sparse array to its pooled coordinates and write into a new hash array to reflect the downsampled sparsity pattern. We then loop over pooled pixel positions and accumulate features from pixels that land in the same pool, finally applying the pooling operation. The algorithmic details are given in Alg. 4.
3.2.5. Sparse flattening
All sparse layers above compute on and output sparse arrays that represent a reduced representation of the original image. Before connecting to conventional dense layers for downstream tasks, we transform back to a flat dense representation. We first initialize a flat dense array with zeros, then map the coordinate of each retained pixel to the corresponding index and scatter its channel features to the corresponding slots. The algorithmic details are given in Alg 5.
4. Evaluations
We perform scaling studies for the five sparse layers to understand their latency and resource utilization on an FPGA by scanning key hyperparameters, as presented in Sec. 4.1. We then perform experiments to evaluate sparse CNNs on an FPGA comparing against standard CNN baselines on three datasets in Sec. 4.2: the datasets are described in Sec. 4.2.1, the experimental setup in Sec. 4.2.2, and the results in Sec. 4.2.3.
4.1. Scaling studies
In the scaling studies, layers are synthesized with Vitis HLS (2023.1) (AMD, 2023), targeting an AMD/Xilinx Alveo FPGA with part number xcu250-figd2104-2L-e. We report FPGA resource utilization from the HLS compilation step (C-synthesis results), which typically provides conservative estimates (overestimation). We quote C-synthesis numbers here rather than logic synthesis because we are characterizing individual layers: logic synthesis can allocate a non-negligible fraction of resources into shared FIFO buffers across layers, which could lead to double-counting if quoted per layer. We report latency in clock cycles, using a 200 MHz clock (a standard benchmark for and deployments at the LHC), and utilization for four resource types: block random access memory (BRAM), digital signal processor (DSP), flip-flop (FF), and look-up table (LUT). Utilization is given as a percentage relative to the total available on the target device: BRAMs5,376, DSPs12,288, FFs3,456,000, and LUTs1,728,000.
For the sparse input reduction layer, we test six input sizes: 500, 1000, 1500, 2000, 2500, and 3000 pixels. For each input size, we scan from 5 to 30 in steps of 5. The latency and resource utilization are shown in Fig. 6. Latency scales linearly with across input sizes because the outer loop iterates over the retained active pixels. The latency curves for 500/1000/1500/2000 pixels coincide and differ from 2500/3000, reflecting the stepwise behavior of the reduction tree depth in Eq. 2 (Sec. 3.2.1). No DSPs are required because the readout involves only comparisons and data moves. FF/LUT usage is nearly flat versus for a fixed input size, dominated by the fixed reduction-tree width, and increases with input size or bit width.
For the sparse convolutional layer, we evaluate , kernel sizes , and from 5 to 30 in steps of 5. We choose a small number of input/output channels because spatially sparse inputs typically exhibit simpler semantic patterns, so compact models suffice for low-latency applications and can achieve performance comparable to standard CNNs (see Sec. 4.2.3). In addition, fully parallel implementations with many filters often incur prohibitive resource usage on an FPGA. Exploring larger numbers of filters is more appropriate once the framework is extended with a streaming implementation in the future. Latencies are shown in Fig. 7 and resource utilization in Fig. 8. Full or partial parallelization of standard convolution at these input sizes is beyond the capabilities of the current implementation, so we do not include it. Latency and resource usage are broadly similar for and , except that the latter can require more BRAMs to store larger weight tensors as channels increase. This matches the design in Sec. 3.2.2: the number of iterations does not scale with because kernel positions are accessed via offset checks rather than explicit loops over all positions. In our tested configurations, the total number of iterations is small enough to fully unroll the loops, yielding very low latency at the cost of higher resource usage. As , , or grows, the degree of parallelism increases and so do resources.
4.2. Evaluation on relevant datasets
4.2.1. Datasets
We perform experiments on three datasets with sparse image data: handwritten digits, neutrino detection in the MicroBooNE experiment, and jet flavor tagging at the LHC. The datasets and preprocessing are described below.
We first inspect a toy dataset derived from MNIST (LeCun and Cortes, 2010). The original images have pixels and are already somewhat sparse, but we target sparsity at the few-percent level or below. For demonstration, we further sparsify the features with the following transformations: (1) apply average pooling with pool size 3 to reduce the number of nonzeros; (2) zero-pad the boarders to resize to ; (3) radially “inflate” pixels outward proportional to their distance from the image center, so the active pixels are more separated spatially (this avoids over-compression by later pooling); (4) zero out low-valued pixels with a threshold at 0.4. An example image is shown in Fig. 9. The task we consider is supervised multi-class classification of the processed images. We also illustrate the importance of choosing a proper , a hyperparameter of our sparse CNNs (and a fixed bound in HLS). As described in Sec. 3.2.1, caps the number of active pixels retained in sparse arrays which are subsequently processed by sparse convolution. Pixels are scanned in a row-major order (top left to bottom right). A larger preserves more information but increases hardware cost, whereas too small a value leads to semantic information loss in a way that depends on the spatial sparsity of a given dataset. In Fig. 10, using the digit “9” as an example, is insufficient to resemble any class; 8 captures only the upper half which resembles the digit “0”; 12 begins to include the lower part; and 16 recovers the full “9”. Thus an scan is necessary to balance accuracy and hardware cost for a given dataset.
Next, we consider a realistic neutrino dataset from MicroBooNE Open Samples (Acciarri et al., 2017b; Abratenko et al., 2022b, a). MicroBooNE uses a liquid argon time projection chamber (LArTPC) to record neutrino interaction events. In a LArTPC, an incoming neutrino (or a cosmic-ray particle) interacts with an argon nucleus, producing charged particles that ionize argon along their trajectories. Ionization electrons drift to the anode under an applied electric field and induce signals on wire sensor planes. The arrival time and wire position are used to reconstruct 2D images containing particle trajectories. The dataset we use contains simulated neutrino interactions overlaid on real cosmic-ray background. Originally, each LArTPC image has 6400 time bins and 3456 wire bins (or 2400, depending on plane), and pixel values are sensor amplitudes. An example is shown at the top of Fig. 11, where the time axis is compressed by a factor of 6, and the pixel intensity is saturated at 100 and cut at 10. The objective is to detect whether an image contains a neutrino interaction. We extract windows of size , which are large enough to contain nearly all of the activity from a neutrino interaction. Windows containing a simulated neutrino interaction (based on truth labels) are labeled as signal, while background windows are randomly cropped elsewhere. To make the computation manageable, we downsample the windows to by summing the bins, then denoise by zeroing pixels below 700, as shown in Fig. 11. Signal windows tend to contain intersecting tracks forming a vertex while background windows contain fewer active pixels or disconnected tracks with low intensity. Finally, features are normalized by a factor of 3200 to unit order. The task we consider is supervised binary classification.
Lastly, we consider jet flavor tagging at the LHC (Duarte et al., 2018; Pierini et al., 2020; Moreno et al., 2020). In proton-proton collisions, unstable heavy particles are produced and they decay promptly into sprays of particles recorded by the detector. When the parent particle has high momentum, the constituent particles appear as a collimated cone, called a jet. Jet substructure (particle content, energy, and geometry) encodes the identity of the parent particle, and identifying the originating particle is called jet tagging. The dataset we use contains simulated jets with transverse momentum () around 1 TeV produced from proton-proton collisions at the center-of-mass energy of 13 TeV. Jets are clustered with the anti- algorithm (Cacciari et al., 2008) using the distance parameter at . Five jet classes are simulated: gloun (g), light quark (q), W boson, Z boson, and top quark (t). Each jet is represented as a image in pseudorapidity (transformed polar angle) and azimuthal angle relative to the jet axis; each pixel stores the scalar sum of particle measured in calorimeters, and up to the 150 highest- constituents per jet are retained. Example images are shown in the top row of Fig. 12. The task we consider is supervised multi-class classification of the processed images. Because the original images are too large for full parallelization on an FPGA, we preprocess the images for demonstration: (1) crop from to by removing boarder pixels; (2) normalize pixel values by a factor of 1200, saturate at 1, and cut at 0.003 (middle row in Fig. 12); (3) as with MNIST, apply radial inflation to spread active pixels outward from the center so that subsequent pooling does not collapse information (bottom row in Fig. 12).
4.2.2. Experimental setup
For a fair comparison, we use the same base architecture to construct both the standard CNN and the sparse CNN in each experiment. The three models, one for each dataset experiment, are shown in Fig. 13. Each model consists of two convolution blocks (a convolutional layer followed by ReLU), uses average pooling for downsampling, then flattens to feed a two-layer multilayer perceptron (MLP) classifier. These models are compact (approx. 4k parameters), as is typical for low-latency FPGA applications. The only architectural difference between the standard and sparse models is that the sparse CNN computes on at most pixels. For the sparse CNNs, we test , referred to as sparse-{tiny,small,medium,large}. We compare accuracy and FPGA cost between the standard CNN and these sparse variants, training with our library using for quantization-aware training, evaluating two fixed total bit-widths: 8 and 16.
For MNIST, we use the 70k examples split into 50k/10k/10k for train/val/test. Models are trained with batch size 128 for up to 200 epochs with early stopping, using Adam (Kingma and Ba, 2017) with a learning rate of to optimize categorical cross-entropy. For MicroBooNE dataset, we use 160k windows with balanced signal and background, split 70/10/20% for train/val/test. Training settings are the same as for MNIST except the loss is binary cross-entropy. For the LHC jet dataset, we use 300k jets split 70/10/20% for train/val/test sets, with the same training settings as for MNIST.
After training, we generate HLS for all models with (Duarte et al., 2018; FastML Team, 2025; Aarrestad et al., 2021). Standard CNNs are implemented with stream I/O (they are too large with parallel I/O) and synthesized with Vivado HLS (2020.1) (Xilinx, 2020), the currently stable version for convolutional architectures. Sparse CNNs use parallel I/O and are synthesized with Vitis HLS (2023.1) (AMD, 2023). The reported resources are obtained after the logic synthesis step. Same as in the scaling studies, the target device is an AMD/Xilinx Alveo FPGA with part number xcu250-figd2104-2L-e, and the clock frequency is set to 200 MHz (5 ns clock period).
4.2.3. Results
Fig. 14 illustrates sparse convolution on an example neutrino signal image from the MicroBooNE dataset, showing the input and the output after one sparse convolutional layer. As designed, the active pixel locations are unchanged. Sparse convolution is a constrained operation, where feature updates occur only at active pixels. This limits the spread of information to neighboring locations and narrows how pixels communicate across layers. Since the number of active pixels never increases (sparse convolution preserves the active set and pooling typically reduces it), one can trade computation for accuracy by increasing the number of output channels per layer. Likewise, a larger kernel can increase the receptive field among active pixels, as our HLS implementation does not directly scale the iteration count with the kernel size (see Sec. 3.2.2).
Fig. 15 shows model performance. As expected, sparse CNN accuracy improves with larger because more semantic information is retained. The sparse-large models () are the most accurate among the sparse variants and are close to the standard CNNs. A small performance loss is expected, since sparse CNNs compute on a subset of pixels and enforce a sparsity-preserving convolution, but these small losses are traded for singificant FPGA inference speedups.
Tab. 1 summarizes latency, speedup, and performance relative to the standard CNNs. Across all datasets, sparse-large gains more than an order-of-magnitude speedup with comparable or slightly lower accuracy. The high latency of standard CNNs comes from computing on all pixels and, given the large image sizes, the need to use stream I/O (one pixel streaming per cycle plus additional buffering). In contrast, sparse CNNs selectively compute on a small fraction of pixels, for example, on MicroBooNE, sparse-large consumes 20 of input pixels (0.5%), and the compact representation enables parallel processing on FPGA, where all active pixels can be processed concurrently. Concretely, for MicroBooNE, sparse-large at 8-bit achieves a speedup over the 8-bit standard CNN (133 vs. 9733 clock cycles, i.e., 0.665 vs. 48.665 s) with % ROC AUC drop (0.927 vs. 0.943). Similar results are observed for the other datasets.
| Sparse MNIST, 10 classes, input pixels | ||||||||||||||
| CNN model | Latency [cc] | Speedup | II [cc] | Acc. [%] | AUC(0) | AUC(1) | AUC(2) | AUC(3) | AUC(4) | AUC(5) | AUC(6) | AUC(7) | AUC(8) | AUC(9) |
| Sparse-t 8b | 79 (0.395 s) | 35 (0.175 s) | 74.3 | 0.9631 | 0.9972 | 0.9562 | 0.9494 | 0.9822 | 0.9666 | 0.9893 | 0.9583 | 0.9038 | 0.9581 | |
| Sparse-t 16b | 81 (0.405 s) | 35 (0.175 s) | 83.1 | 0.9834 | 0.9990 | 0.9788 | 0.9735 | 0.9941 | 0.9872 | 0.9948 | 0.9817 | 0.9476 | 0.9770 | |
| Sparse-s 8b | 104 (0.52 s) | 52 (0.26 s) | 86.7 | 0.9967 | 0.9981 | 0.9886 | 0.9839 | 0.9906 | 0.9920 | 0.9959 | 0.9859 | 0.9762 | 0.9813 | |
| Sparse-s 16b | 106 (0.53 s) | 52 (0.26 s) | 91.8 | 0.9987 | 0.9994 | 0.9956 | 0.9922 | 0.9969 | 0.9971 | 0.9982 | 0.9943 | 0.9893 | 0.9908 | |
| Sparse-m 8b | 124 (0.62 s) | 67 (0.335 s) | 90.7 | 0.9984 | 0.9987 | 0.9956 | 0.9914 | 0.9942 | 0.9946 | 0.9980 | 0.9919 | 0.9892 | 0.9882 | |
| Sparse-m 16b | 127 (0.635 s) | 67 (0.335 s) | 94.6 | 0.9994 | 0.9996 | 0.9984 | 0.9963 | 0.9978 | 0.9983 | 0.9991 | 0.9975 | 0.9966 | 0.9940 | |
| Sparse-l 8b | 146 (0.73 s) | 84 (0.42 s) | 91.2 | 0.9988 | 0.9982 | 0.9957 | 0.9925 | 0.9942 | 0.9955 | 0.9984 | 0.9929 | 0.9906 | 0.9899 | |
| Sparse-l 16b | 150 (0.75 s) | 84 (0.42 s) | 95.0 | 0.9991 | 0.9997 | 0.9989 | 0.9969 | 0.9987 | 0.9979 | 0.9991 | 0.9971 | 0.9974 | 0.9952 | |
| Standard 8b | 3232 (16.16 s) | 3124 (15.62 s) | 94.3 | 0.9991 | 0.9997 | 0.9982 | 0.9962 | 0.9980 | 0.9977 | 0.9985 | 0.9968 | 0.9958 | 0.9961 | |
| Standard 16b | 3232 (16.16 s) | 3124 (15.62 s) | 95.4 | 0.9996 | 0.9997 | 0.9989 | 0.9978 | 0.9988 | 0.9981 | 0.9989 | 0.9985 | 0.9979 | 0.9976 | |
| MicroBooNE neutrino detection, 2 classes, input pixels | |||||
| CNN model | Latency [cc] | Speedup | II [cc] | AUC | |
| Sparse-t 8b | 69 (0.345 s) | 35 (0.175 s) | 0.819 | 0.887 | |
| Sparse-t 16b | 71 (0.355 s) | 35 (0.175 s) | 0.830 | 0.891 | |
| Sparse-s 8b | 90 (0.45 s) | 52 (0.26 s) | 0.856 | 0.900 | |
| Sparse-s 16b | 94 (0.47 s) | 52 (0.26 s) | 0.864 | 0.908 | |
| Sparse-m 8b | 111 (0.555 s) | 67 (0.335 s) | 0.879 | 0.913 | |
| Sparse-m 16b | 116 (0.58 s) | 67 (0.335 s) | 0.881 | 0.918 | |
| Sparse-l 8b | 133 (0.665 s) | 84 (0.42 s) | 0.896 | 0.927 | |
| Sparse-l 16b | 136 (0.68 s) | 84 (0.42 s) | 0.894 | 0.924 | |
| Standard 8b | 9733 (48.665 s) | 5029 (25.145 s) | 0.921 | 0.943 | |
| Standard 16b | 9736 (48.68 s) | 5029 (25.145 s) | 0.921 | 0.940 | |
| LHC jet tagging, 5 classes, input pixels | |||||||||
| CNN model | Latency [cc] | Speedup | II [cc] | Acc. [%] | AUC(g) | AUC(q) | AUC(W) | AUC(Z) | AUC(t) |
| Sparse-t 8b | 75 (0.375 s) | 35 (0.175 s) | 53.0 | 0.7853 | 0.8443 | 0.8122 | 0.7858 | 0.8787 | |
| Sparse-t 16b | 78 (0.39 s) | 35 (0.175 s) | 55.8 | 0.7976 | 0.8494 | 0.8388 | 0.8117 | 0.8822 | |
| Sparse-s 8b | 97 (0.485 s) | 52 (0.26 s) | 59.0 | 0.8195 | 0.8626 | 0.8563 | 0.8328 | 0.8996 | |
| Sparse-s 16b | 101 (0.505 s) | 52 (0.26 s) | 61.5 | 0.8340 | 0.8685 | 0.8779 | 0.8529 | 0.9053 | |
| Sparse-m 8b | 119 (0.595 s) | 67 (0.335 s) | 61.8 | 0.8418 | 0.8722 | 0.8866 | 0.8619 | 0.9130 | |
| Sparse-m 16b | 120 (0.6 s) | 67 (0.335 s) | 65.5 | 0.8580 | 0.8786 | 0.9029 | 0.8856 | 0.9238 | |
| Sparse-l 8b | 140 (0.7 s) | 84 (0.42 s) | 65.7 | 0.8617 | 0.8805 | 0.9077 | 0.8856 | 0.9238 | |
| Sparse-l 16b | 143 (0.715 s) | 84 (0.42 s) | 68.1 | 0.8710 | 0.8852 | 0.9208 | 0.9032 | 0.9328 | |
| Standard 8b | 8390 (41.95 s) | 4340 (21.7 s) | 68.3 | 0.8468 | 0.8743 | 0.9324 | 0.9178 | 0.9374 | |
| Standard 16b | 8390 (41.95 s) | 4340 (21.7 s) | 72.2 | 0.8874 | 0.8893 | 0.9463 | 0.9361 | 0.9426 | |
For all sparse CNNs, the initiation interval (II)—the number of clock cycles between accepting successive inputs—equals the latency of the sparse input reduction layer (the first layer in sparse models), which can be seen in the per-layer latency breakdown in Fig. 17 (8-bit) and Fig. 22 (16-bit). This is because the recursive tree splitting requires buffering and processing the current input array until scanning completes before the next array can enter. Subsequent sparse layers operate on sparse arrays and their operations are maximally parallelized.
Fig. 16 (Fig. 21) shows resource usage as a fraction of the total available on device for 8-bit (16-bit) models. As discussed, standard CNNs, implemented with stream I/O, tend to use fewer resources due to sequential processing, while sparse CNNs, implemented with parallel I/O, consume more resources in exchange for much lower latency. For example, LUT usage is several times higher in sparse CNNs because loops are largely unrolled to execute concurrently. This increase is less concerning since even the sparse-large models consume only of total LUTs on device, but this is traded for much larger speedups. When needed, reuse factors—an HLS setting that controls parallelization by having a processing element perform more sequential operations—can be increased for the sparse CNNs to reduce resources at the cost of higher latency, a trade-off that is acceptable given the currently seen substantial baseline speedups. Per-layer resource breakdowns shown in Fig. 17 Fig. 22 are consistent with the scaling behaviors in Sec. 4.1.
5. Summary
We have introduced , a framework for sub-microsecond FPGA inference of CNNs using sparse convolution on spatially sparse 2D data. The key features in our HLS implementation are: (1) dynamically identifying and packing active pixels, a small subset of a large input, into compact arrays, then (2) performing all subsequent computation only on those packed pixels while ignoring inactive pixels. By selectively computing on the necessarily informative subset, the design avoids computation on empty regions, yielding significant inference speedups and resource savings, especially when only percent-level or fewer pixels are active. As a representative example on MicroBooNE neutrino images with 4000 input pixels, a sparse CNN that computes on at most 20 active pixels achieves a inference speedup on an FPGA over a standard CNN with the same base architecture (133 vs. 9733 clock cycles; 0.665 vs. 48.665 s), with ROC AUC drop, and resource utilization well within device budgets (e.g., on the order of 20% LUTs for the largest sparse configuration). For easy adoption, we provide a Python library supporting quantization-aware training of sparse CNNs and HLS C++ modules for FPGA deployment. Codes are available at https://github.com/hftsoi/sparse-pixels, and automated HLS generation will be supported in a future version of . Future directions include implementing sparse layers with stream I/O, exploring partial parallelization, and integrating the framework with high-granularity quantization. We hope this work benefits experiments operating under stringent sub-microsecond latency constraints, such as modern particle experiments, where standard CNNs are often infeasible due to too high a latency and would otherwise force substantial performance compromises.
Acknowledgements.
DR is supported by the U.S. Department of Energy (DOE), Office of Science, Office of High Energy Physics Early Career Research program under Award No. DE-SC0025324. PH is supported by the U.S. National Science Foundation (NSF) Harnessing the Data Revolution (HDR) Institute for Accelerating AI Algorithms for Data Driven Discovery (A3D3) under Cooperative Agreement PHY-2117997. PH is also supported by the Institute for Artificial Intelligence and Fundamental Interactions (IAIFI), under the NSF grant #PHY-2019786. We acknowledge the MicroBooNE Collaboration for making publicly available the data sets (Abratenko et al., 2022b, a) employed in this work. These data sets consist of simulated neutrino interactions from the Booster Neutrino Beamline overlaid on top of cosmic data collected with the MicroBooNE detector (Acciarri et al., 2017b). This work used resources available through the National Research Platform (NRP) at the University of California, San Diego (Weitzel et al., 2025). NRP has been developed, and is supported in part, by funding from National Science Foundation, from awards 1730158, 1540112, 1541349, 1826967, 2112167, 2100237, and 2120019, as well as additional funding from community partners.References
- (1)
- Aad et al. (2020) Georges Aad et al. 2020. Operation of the ATLAS trigger system in Run 2. JINST 15, 10 (2020), P10004. arXiv:2007.12539 [physics.ins-det] doi:10.1088/1748-0221/15/10/P10004
- Aarrestad et al. (2021) Thea Aarrestad et al. 2021. Fast convolutional neural networks on FPGAs with hls4ml. Mach. Learn. Sci. Tech. 2, 4 (2021), 045015. arXiv:2101.05108 [cs.LG] doi:10.1088/2632-2153/ac0ea1
- Aarrestad et al. (2022) Thea Aarrestad et al. 2022. The Dark Machines Anomaly Score Challenge: Benchmark Data and Model Independent Event Classification for the Large Hadron Collider. SciPost Phys. 12, 1 (2022), 043. arXiv:2105.14027 [hep-ph] doi:10.21468/SciPostPhys.12.1.043
- Abi et al. (2020) B. Abi et al. 2020. Neutrino interaction classification with a convolutional neural network in the DUNE far detector. Phys. Rev. D 102, 9 (2020), 092003. arXiv:2006.15052 [physics.ins-det] doi:10.1103/PhysRevD.102.092003
- Abratenko et al. (2022a) Polina Abratenko et al. 2022a. MicroBooNE BNB Electron Neutrino Overlay Sample (With Wire Info). doi:10.5281/zenodo.7262140
- Abratenko et al. (2022b) Polina Abratenko et al. 2022b. MicroBooNE BNB Inclusive Overlay Sample (With Wire Info). doi:10.5281/zenodo.7262009
- Acciarri et al. (2017a) R. Acciarri et al. 2017a. Convolutional Neural Networks Applied to Neutrino Events in a Liquid Argon Time Projection Chamber. JINST 12, 03 (2017), P03011. arXiv:1611.05531 [physics.ins-det] doi:10.1088/1748-0221/12/03/P03011
- Acciarri et al. (2017b) R. Acciarri et al. 2017b. Design and Construction of the MicroBooNE Detector. JINST 12, 02 (2017), P02017. arXiv:1612.05824 [physics.ins-det] doi:10.1088/1748-0221/12/02/P02017
- AMD (2023) AMD. 2023. Vitis Libraries (2023.1). https://docs.amd.com/r/2023.1-English/Vitis_Libraries/index.html.
- Aurisano et al. (2016) A. Aurisano, A. Radovic, D. Rocco, A. Himmel, M. D. Messier, E. Niner, G. Pawloski, F. Psihas, A. Sousa, and P. Vahle. 2016. A Convolutional Neural Network Neutrino Event Classifier. JINST 11, 09 (2016), P09001. arXiv:1604.01444 [hep-ex] doi:10.1088/1748-0221/11/09/P09001
- Blott et al. (2018) Michaela Blott, Thomas B Preußer, Nicholas J Fraser, Giulio Gambardella, Kenneth O’brien, Yaman Umuroglu, Miriam Leeser, and Kees Vissers. 2018. FINN-R: An end-to-end deep-learning framework for fast exploration of quantized neural networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS) 11, 3 (2018), 1–23.
- Butter et al. (2019) Anja Butter et al. 2019. The Machine Learning landscape of top taggers. SciPost Phys. 7 (2019), 014. arXiv:1902.09914 [hep-ph] doi:10.21468/SciPostPhys.7.1.014
- Cacciari et al. (2008) Matteo Cacciari, Gavin P. Salam, and Gregory Soyez. 2008. The anti- jet clustering algorithm. JHEP 04 (2008), 063. arXiv:0802.1189 [hep-ph] doi:10.1088/1126-6708/2008/04/063
- Chollet et al. (2015) François Chollet et al. 2015. Keras. https://keras.io.
- Choy et al. (2019) Christopher Choy, JunYoung Gwak, and Silvio Savarese. 2019. 4D Spatio-Temporal ConvNets: Minkowski Convolutional Neural Networks. In Proceedings of the IEEE Conference on Computer Vision and Pattern Recognition. 3075–3084.
- Chung et al. (2025) Seokju Chung, Jack Cleeve, Akshay Malige, Georgia Karagiorgi, Lino Gerlach, Adrian A. Pol, and Isobel Ojalvo. 2025. Real-time Anomaly Detection for Liquid Argon Time Projection Chambers. (9 2025). arXiv:2509.21817 [hep-ex]
- Coccaro et al. (2023) Andrea Coccaro, Francesco Armando Di Bello, Stefano Giagu, Lucrezia Rambelli, and Nicola Stocchetti. 2023. Fast neural network inference on FPGAs for triggering on long-lived particles at colliders. Mach. Learn. Sci. Tech. 4, 4 (2023), 045040. arXiv:2307.05152 [hep-ex] doi:10.1088/2632-2153/ad087a
- Coelho et al. (2021) Claudionor N. Coelho, Aki Kuusela, Shan Li, Hao Zhuang, Thea Aarrestad, Vladimir Loncar, Jennifer Ngadiuba, Maurizio Pierini, Adrian Alan Pol, and Sioni Summers. 2021. Automatic heterogeneous quantization of deep neural networks for low-latency inference on the edge for particle detectors. Nature Mach. Intell. 3 (2021), 675–686. arXiv:2006.10159 [physics.ins-det] doi:10.1038/s42256-021-00356-5
- Collaboration (2018) ATLAS Collaboration. 2018. Technical Design Report for the Phase-II Upgrade of the ATLAS TDAQ System. ATLAS Technical Design Report CERN-LHCC-2017-020, ATLAS-TDR-029 (2018). doi:10.17181/CERN.2LBB.4IAL
- Collaboration (2020) CMS Collaboration. 2020. The Phase-2 Upgrade of the CMS Level-1 Trigger. CMS Technical Design Report CERN-LHCC-2020-004, CMS-TDR-021 (2020).
- Collaboration (2023) CMS Collaboration. 2023. Level-1 Trigger Calorimeter Image Convolutional Anomaly Detection Algorithm. (2023). https://cds.cern.ch/record/2879816
- Collaboration (2024) CMS Collaboration. 2024. Model-Independent Real-Time Anomaly Detection at the CMS Level-1 Calorimeter Trigger with CICADA. (2024). https://cds.cern.ch/record/2917884
- de Oliveira et al. (2016) Luke de Oliveira, Michael Kagan, Lester Mackey, Benjamin Nachman, and Ariel Schwartzman. 2016. Jet-images — deep learning edition. JHEP 07 (2016), 069. arXiv:1511.05190 [hep-ph] doi:10.1007/JHEP07(2016)069
- Dominé and Terao (2020) Laura Dominé and Kazuhiro Terao. 2020. Scalable deep convolutional neural networks for sparse, locally dense liquid argon time projection chamber data. Phys. Rev. D 102, 1 (2020), 012005. arXiv:1903.05663 [hep-ex] doi:10.1103/PhysRevD.102.012005
- Duarte et al. (2018) Javier Duarte et al. 2018. Fast inference of deep neural networks in FPGAs for particle physics. JINST 13, 07 (2018), P07027. arXiv:1804.06913 [physics.ins-det] doi:10.1088/1748-0221/13/07/P07027
- FastML Team (2025) FastML Team. 2025. fastmachinelearning/hls4ml. doi:10.5281/zenodo.1201549
- Francescato et al. (2021) Simone Francescato et al. 2021. Model compression and simplification pipelines for fast deep neural network inference in FPGAs in HEP. Eur. Phys. J. C 81, 11 (2021), 969. doi:10.1140/epjc/s10052-021-09770-w [Erratum: Eur.Phys.J.C 81, 1064 (2021)].
- Gandrakota (2025) Abhijith Gandrakota. 2025. Realtime Anomaly Detection at the L1 Trigger of CMS Experiment. PoS ICHEP2024 (2025), 1025. arXiv:2411.19506 [hep-ex] doi:10.22323/1.476.1025
- Ghielmetti et al. (2022) Nicolò Ghielmetti et al. 2022. Real-time semantic segmentation on FPGAs for autonomous vehicles with hls4ml. Mach. Learn. Sci. Tech. (2022). arXiv:2205.07690 [cs.CV] doi:10.1088/2632-2153/ac9cb5
- Govorkova et al. (2022) Ekaterina Govorkova et al. 2022. Autoencoders on field-programmable gate arrays for real-time, unsupervised new physics detection at 40 MHz at the Large Hadron Collider. Nature Mach. Intell. 4 (2022), 154–161. arXiv:2108.03986 [physics.ins-det] doi:10.1038/s42256-022-00441-3
- Graham (2015) Ben Graham. 2015. Sparse 3D convolutional neural networks. arXiv:1505.02890 [cs.CV] https://arxiv.org/abs/1505.02890
- Graham et al. (2018) Benjamin Graham, Martin Engelcke, and Laurens van der Maaten. 2018. 3D Semantic Segmentation with Submanifold Sparse Convolutional Networks. In 2018 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). 9224–9232. arXiv:1711.10275 doi:10.1109/CVPR.2018.00961
- Jiang et al. (2025) Zhixing Jiang, Dennis Yin, Yihui Chen, Elham E. Khoda, Scott Hauck, Shih-Chieh Hsu, Ekaterina Govorkova, Philip Harris, Vladimir Loncar, and Eric A. Moreno. 2025. Low latency transformer inference on FPGAs for physics applications with hls4ml. JINST 20, 04 (2025), P04014. arXiv:2409.05207 [cs.LG] doi:10.1088/1748-0221/20/04/P04014
- Kasieczka et al. (2021) Gregor Kasieczka et al. 2021. The LHC Olympics 2020 a community challenge for anomaly detection in high energy physics. Rept. Prog. Phys. 84, 12 (2021), 124201. arXiv:2101.08320 [hep-ph] doi:10.1088/1361-6633/ac36b9
- Kasieczka et al. (2017) Gregor Kasieczka, Tilman Plehn, Michael Russell, and Torben Schell. 2017. Deep-learning Top Taggers or The End of QCD? JHEP 05 (2017), 006. arXiv:1701.08784 [hep-ph] doi:10.1007/JHEP05(2017)006
- Khoda et al. (2023) Elham E. Khoda et al. 2023. Ultra-low latency recurrent neural network inference on FPGAs for physics applications with hls4ml. Mach. Learn. Sci. Tech. 4, 2 (2023), 025004. arXiv:2207.00559 [cs.LG] doi:10.1088/2632-2153/acc0d7
- Kingma and Ba (2017) Diederik P. Kingma and Jimmy Ba. 2017. Adam: A Method for Stochastic Optimization. arXiv:1412.6980 [cs.LG]
- LeCun et al. (2015) Yann LeCun, Yoshua Bengio, and Geoffrey Hinton. 2015. Deep learning. Nature 521 (2015), 436–444. doi:10.1038/nature14539
- LeCun and Cortes (2010) Yann LeCun and Corinna Cortes. 2010. MNIST handwritten digit database. http://yann.lecun.com/exdb/mnist/. (2010). http://yann.lecun.com/exdb/mnist/
- Liu et al. (2015) Baoyuan Liu, Min Wang, Hassan Foroosh, Marshall Tappen, and Marianna Penksy. 2015. Sparse Convolutional Neural Networks. In 2015 IEEE Conference on Computer Vision and Pattern Recognition (CVPR). 806–814. doi:10.1109/CVPR.2015.7298681
- Lu et al. (2019) Liqiang Lu, Jiaming Xie, Ruirui Huang, Jiansong Zhang, Wei Lin, and Yun Liang. 2019. An Efficient Hardware Accelerator for Sparse Convolutional Neural Networks on FPGAs. In 2019 IEEE 27th Annual International Symposium on Field-Programmable Custom Computing Machines (FCCM). 17–25. doi:10.1109/FCCM.2019.00013
- Maglianella et al. (2023) Leandro Maglianella, Lorenzo Nicoletti, Stefano Giagu, Christian Napoli, and Simone Scardapane. 2023. Convergent Approaches to AI Explainability for HEP Muonic Particles Pattern Recognition. Comput. Softw. Big Sci. 7, 1 (2023), 8. doi:10.1007/s41781-023-00102-z
- Moreno et al. (2020) Eric A. Moreno, Olmo Cerri, Javier M. Duarte, Harvey B. Newman, Thong Q. Nguyen, Avikar Periwal, Maurizio Pierini, Aidana Serikova, Maria Spiropulu, and Jean-Roch Vlimant. 2020. JEDI-net: a jet identification algorithm based on interaction networks. Eur. Phys. J. C 80, 1 (2020), 58. arXiv:1908.05318 [hep-ex] doi:10.1140/epjc/s10052-020-7608-4
- Paszke et al. (2019) Adam Paszke, Sam Gross, Francisco Massa, Gal Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Naresh Gimelshein, Luca Antiga, et al. 2019. PyTorch: An Imperative Style, High-Performance Deep Learning Library. Advances in Neural Information Processing Systems 32 (2019).
- Pierini et al. (2020) Maurizio Pierini, Javier Mauricio Duarte, Nhan Tran, and Marat Freytsis. 2020. HLS4ML LHC Jet dataset (150 particles). doi:10.5281/zenodo.3602260
- Pol et al. (2021) Adrian Alan Pol et al. 2021. Jet Single Shot Detection. EPJ Web Conf. 251 (2021), 04027. arXiv:2105.05785 [hep-ex] doi:10.1051/epjconf/202125104027
- Sirunyan et al. (2020) Albert M Sirunyan et al. 2020. Performance of the CMS Level-1 trigger in proton-proton collisions at 13 TeV. JINST 15, 10 (2020), P10017. arXiv:2006.10165 [hep-ex] doi:10.1088/1748-0221/15/10/P10017
- Summers et al. (2020) Sioni Summers et al. 2020. Fast inference of Boosted Decision Trees in FPGAs for particle physics. JINST 15, 05 (2020), P05026. arXiv:2002.02534 [physics.comp-ph] doi:10.1088/1748-0221/15/05/P05026
- Sun et al. (2024) Chang Sun, Thea K. Arrestad, Vladimir Loncar, Jennifer Ngadiuba, and Maria Spiropulu. 2024. Gradient-based Automatic Mixed Precision Quantization for Neural Networks On-Chip. (5 2024). arXiv:2405.00645 [cs.LG] doi:10.7907/hq8jd-rhg30
- Tang et al. (2022) Haotian Tang, Zhijian Liu, Xiuyu Li, Yujun Lin, and Song Han. 2022. TorchSparse: Efficient Point Cloud Inference Engine. In Conference on Machine Learning and Systems (MLSys).
- Tang et al. (2023) Haotian Tang, Shang Yang, Zhijian Liu, Ke Hong, Zhongming Yu, Xiuyu Li, Guohao Dai, Yu Wang, and Song Han. 2023. TorchSparse++: Efficient Training and Inference Framework for Sparse Convolution on GPUs. In IEEE/ACM International Symposium on Microarchitecture (MICRO).
- Tarafdar et al. (2022) Naif Tarafdar, Giuseppe Di Guglielmo, Philip C. Harris, Jeffrey D. Krupa, Vladimir Loncar, Dylan S. Rankin, Nhan Tran, Zhenbin Wu, Qianfeng Clark Shen, and Paul Chow. 2022. AIgean: An Open Framework for Deploying Machine Learning on Heterogeneous Clusters. ACM Trans. Reconf. Tech. Syst. 15, 3 (2022), 1–32. doi:10.1145/3482854
- Tsoi et al. (2025) Ho Fung Tsoi, Vladimir Loncar, Sridhara Dasu, and Philip Harris. 2025. SymbolNet: neural symbolic regression with adaptive dynamic pruning for compression. Mach. Learn. Sci. Tech. 6, 1 (2025), 015021. arXiv:2401.09949 [cs.LG] doi:10.1088/2632-2153/adaad8
- Tsoi et al. (2024) Ho Fung Tsoi, Adrian Alan Pol, Vladimir Loncar, Ekaterina Govorkova, Miles Cranmer, Sridhara Dasu, Peter Elmer, Philip Harris, Isobel Ojalvo, and Maurizio Pierini. 2024. Symbolic Regression on FPGAs for Fast Machine Learning Inference. EPJ Web Conf. 295 (2024), 09036. arXiv:2305.04099 [cs.LG] doi:10.1051/epjconf/202429509036
- Umuroglu et al. (2017) Yaman Umuroglu, Nicholas J. Fraser, Giulio Gambardella, Michaela Blott, Philip Leong, Magnus Jahre, and Kees Vissers. 2017. FINN: A Framework for Fast, Scalable Binarized Neural Network Inference. In Proceedings of the 2017 ACM/SIGDA International Symposium on Field-Programmable Gate Arrays (FPGA ’17). ACM, 65–74.
- Vreča and Biasizzo (2025) Jure Vreča and Anton Biasizzo. 2025. Generating Direct Logic Circuit Implementations of Deeply Quantized Neural Networks Using Chisel4ml. Electronics 14, 5 (2025). doi:10.3390/electronics14050849
- Weitzel et al. (2025) Derek Weitzel, Ashton Graves, Sam Albin, Huijun Zhu, Frank Wuerthwein, Mahidhar Tatineni, Dmitry Mishin, Elham Khoda, Mohammad Sada, Larry Smarr, Thomas DeFanti, and John Graham. 2025. The National Research Platform: Stretched, Multi-Tenant, Scientific Kubernetes Cluster. In Practice and Experience in Advanced Research Computing 2025: The Power of Collaboration (PEARC ’25). Article 69, 5 pages. doi:10.1145/3708035.3736060
- Xilinx (2020) Xilinx. 2020. Vivado Design Suite User Guide: High-Level Synthesis. https://www.xilinx.com/support/documentation/sw_manuals/xilinx2020_1/ug902-vivado-high-level-synthesis.pdf.
- Yan et al. (2018) Yan Yan, Yuxing Mao, and Bo Li. 2018. SECOND: Sparsely Embedded Convolutional Detection. Sensors 18, 10 (2018). doi:10.3390/s18103337
- Zhu et al. (2020) Chaoyang Zhu, Kejie Huang, Shuyuan Yang, Ziqi Zhu, Hejia Zhang, and Haibin Shen. 2020. An Efficient Hardware Accelerator for Structured Sparse Convolutional Neural Networks on FPGAs. IEEE Trans. Very Large Scale Integr. Syst. 28, 9 (Sept. 2020), 1953–1965. doi:10.1109/TVLSI.2020.3002779
- Zurbano Fernandez et al. (2020) I. Zurbano Fernandez et al. 2020. High-Luminosity Large Hadron Collider (HL-LHC): Technical design report. 10/2020 (12 2020). doi:10.23731/CYRM-2020-0010
Appendix A appendix
A.1. Sparse activation in HLS and scaling
The algorithmic details are given in Alg. 3. For scaling, we use ReLU and scan from 5 to 30 in steps of 5. Fig. 18 reports latency and resource utilization. The activation loop is fully unrolled, so the layer finishes in one clock cycle and uses no DSPs. FF/LUT usage increases with because more elements are processed in parallel.
A.2. Sparse pooling in HLS and scaling
The algorithmic details are given in Alg. 4. For scaling, we use average-pooling as an example, test , and scan from 5 to 30 in steps of 5. Fig. 19 shows latency and resource utilization. The latency curves across coincide because the channel loop is unrolled. No DSPs are used; FF/LUT usage increases with and due to more parallel comparisons and accumulators.
A.3. Sparse flatten in HLS and scaling
The algorithmic details are given in Alg. 5. For scaling, we test flat output sizes and scan from 5 to 30 in steps of 5. Fig. 20 shows latency and resource utilization. The latency is essentially independent of the small flat output size because initialization is fully unrolled. Because this is address generation plus writes, DSP usage is 0, while FF/LUT usage increases with and .
A.4. Synthesis results for 16-bit models
Fig. 21 and Fig. 22 summarize results of the 16-bit models. Latencies are similar to the 8-bit models because of same loop structures in architectural design. DSP usage increases primarily because arithmetic such as multiplications are preferably mapped to DSPs at higher bit-width by Vivado.

### v2

@@
 Abstract.
-Inference of standard convolutional neural networks (CNNs) on Field-Programmable Gate Arrays (FPGAs) often incurs high latency and a long initiation interval due to the deep nested loops required to densely convolve every input pixel regardless of its feature value, especially when the image size is large.
-However, in some image data, input features can be spatially sparse, and semantic information may occupy only a small fraction of the input pixels.
-In this case most computation would be wasted on empty regions.
-In this work, we introduce SparsePixels, a framework for efficient convolution for spatially sparse image data on FPGAs, targeting fast inference applications in constrained environments with latency requirements of microseconds or below.
-Our approach implements a special class of CNNs that selectively retain and compute on a small subset of pixels that are active while ignoring the rest.
-We show that, for example, in a neutrino physics dataset for identifying neutrino interactions in LArTPC images that have around 4k input pixels but are naturally very sparse, a standard CNN with a compact size of 4k parameters incurs an inference latency of 48.665 s on an FPGA, whereas a sparse CNN of the same base architecture computing on less than 1% of the input pixels results in a inference speedup to 0.665 s, with resource utilization well within on-chip budgets, trading only a small percent-level performance loss.
-At least one-order-of magnitude speedups with comparable performance are also demonstrated in similar datasets with sparse image patterns.
-This work aims to benefit future algorithm developments for fast and efficient data readout in modern experiments such as the trigger and data acquisition systems at the CERN Large Hadron Collider.
-For easy adoption, we have developed a library to support building sparse CNNs with quantization-aware training, as well as an HLS implementation for FPGA deployment.
-Codes are available at https://github.com/hftsoi/sparse-pixels.
+Inference of standard convolutional neural networks (CNNs) on FPGAs often incurs high latency and a long initiation interval due to the deep nested loops required to densely convolve every input pixel regardless of its feature value.
+However, input features can be spatially sparse in some image data, where semantic information may occupy only a small fraction of the pixels and most computation would be wasted on empty regions.
+In this work, we introduce SparsePixels, a framework that implements sparse convolution on FPGAs by selectively retaining and computing on a small subset of active pixels while ignoring the rest.
+We show that, for identifying neutrino interactions in naturally sparse LArTPC images with 4k pixels, a standard CNN with a compact size of 4k parameters incurs an inference latency of 48.665 s on an FPGA, whereas a sparse CNN of the same base architecture, computing on less than 1% of the input pixels, achieves a speedup to 0.665 s with resource utilization well within on-chip budgets, trading only a small percent-level performance loss.
+This work aims to benefit future algorithm development for efficient data readout in modern experiments with latency requirements of microseconds or below.
+Code is publicly released at https://github.com/hftsoi/sparse-pixels.
 1. Introduction
@@
 Development of a sparse convolution method designed for low-latency FPGA implementation.
-Proof-of-concept on spatially sparse physics image data, demonstrating significant FPGA inference speedups with comparable model performance.
-A Python library for quantization-aware training of sparse CNNs and C++ HLS modules for FPGA deployment.
+Proof-of-concept on spatially sparse image data, demonstrating significant FPGA inference speedups with comparable model performance.
+A Python library for quantization-aware training of sparse CNNs and C++ HLS modules for FPGA deployment. Code is publicly released at https://github.com/hftsoi/sparse-pixels.
 2. Related work
@@
 For easy adoption, we provide a Python library supporting quantization-aware training of sparse CNNs and HLS C++ modules for FPGA deployment.
-Codes are available at https://github.com/hftsoi/sparse-pixels, and automated HLS generation will be supported in a future version of .
 Future directions include implementing sparse layers with stream I/O, exploring partial parallelization, and integrating the framework with high-granularity quantization.